Use the principle of induction Prove that 2 1 n 2 n 2
n n 1
> + ∀ >

,
1
Expert's answer
2019-09-20T11:14:23-0400

Prove that "2^n\geq2n, \forall n\geq1"

For any integer n ≥ 1, let "P_n" be the statement that


"2^n\geq2n"

Base Case. The statement "P_1" says that


"2^1\geq2\cdot1"

which is true.

Inductive Step. Fix "k\geq1," and suppose that "P_k" holds, that is,


"2^k\geq2k"

It remains to show that "P_{k+1}" holds, that is,


"2^{k+1}\geq2(k+1)"


"2k\geq2, \forall k\geq1""2^k\geq2k=>2\cdot2^k\geq2\cdot2k=>""=>2^{k+1}\geq4k=2k+2k\geq2k+2=2(k+1)"

Therefore "P_{k+1}" holds.

Thus, by the principle of mathematical induction, for all "n\geq1,P_n" holds.